OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (2, 7) and (-2, 9).

OpenStudy (anonymous):

Not looking for an easy answer, if possible could you explain to me how to work through this problem?

OpenStudy (littlebird):

Slope intercept form: y=mx+b

OpenStudy (anonymous):

first, find the slope of the line: clue: use the slope formula second, plug the slope in y=mx+b (slope intercept form) third, using the given point find the y-intercept of the equation (b) lastly, rewrite the slope(m) and y-int(b) into y=mx+b (slope intercept form).

OpenStudy (littlebird):

\[\frac{ y1-y2 }{ x1-x2 }\] Slope formula

OpenStudy (anonymous):

I got y=1/2x+8, is that right?

OpenStudy (wolf1728):

(2,7) (-2,9) slope = m = (7 -9) / (2 --2) = -2 / 4 = -1 / 2 Forming the equation b = y -mx b = y -1/2 * x Now take one of the given points (2,7) and plug it into the equation That should give you a darned good start.

OpenStudy (wolf1728):

You know how to form an equation. y = -1/2 x + 8 is right :-)